Juja farm company has an ageing piece of equipment which is less efficient that more modern equivalents. This equipment will continue to operate or another 15 years but operating and maintenance costs will be Shs 3,500,000 per year. Alternatively it could be sold, raising shs. 2,000,000 more and replaced with modern equivalent which costs Shs 7,000,000 but has reduced operating maintenance costs at Shs 3,000,000 per year. This machine could be sold at the end of its 15 yea life for scrap for Shs 500,000. The third possibility is to spend Shs. 2,500,000 for an immediate overhaul of the old machine which will improve its efficiency for the rest of its life. So that operating and maintenance costs become Sh 3,200,000 per annum. The old equipment will have a zero scrap value in 15 years whether or not it is overhauled. Juja farm requires a return of 9% on projects in this risk class.
1
Expert's answer
2021-08-23T13:46:24-0400

In this, we have considered the present value of all methods and select the best among them.

First option

Interest rate=9%

Remaining life=15 years

Annual maintenance cost=3500000

Present value factor"=\frac{1−(1+r)

^{−n}}{

r}

=

\frac{1-(1.09)^{-nr}}{(0.09)}

=8.0606884299"

Present value of cost= Annual cost x present value factor "=3500000 * 8.0606884299"

Net Present Value of option1=-Shs 28212409.504

Investment=Cost of modern equipment - selling price of old equipment "=7000000-2000000=\$5000000"

Annual saving =30000000 per year

Present value of annual savings "=8.0606884299 * 3000000=\$24182065.29"

Present value of salvage value

"\frac{Scrap \space value}{(1+r)^{
n}}\\

=\frac{500000}{(1.09)^{15}}=\$137269.02"

Net present value of second option= Present value of the salvage value